# Cohort-Based Active Modality Acquisition

### v2

@@ C.2 Oracle Acquisition Strategies: Exact Gain Calculation
Oracle acquisition strategies serve as theoretical upper limits for the performance of greedy acquisition approaches. They operate under the ideal assumption that the true labels ${\textnormal{y}}_{i}$ and the outcome scores ${\textnormal{s}}_{i}^{\textnormal{full}}$ are known for all samples $i\in\{1,\dots,N\}$. While not implementable in practice, these oracle strategies provide benchmarks by selecting samples based on their exact marginal contribution to the global evaluation metric. The general principle is to iteratively select $\beta$ samples. At each step, among the samples for which the additional modality has not yet been acquired, the oracle picks the one that provides the largest true immediate gain to the chosen global metric.
@@ C.3 Upper-Bound Heuristic Strategies
The preceding oracle strategies make the assumption of perfect foresight into both the true labels ${\textnormal{y}}_{i}$ and the exact outcome scores ${\textnormal{s}}_{i}^{\textnormal{full}}$. We now introduce a distinct class of upper-bound heuristic strategies. These strategies still presume access to the true future scores ${\textnormal{s}}_{i}^{\textnormal{full}}$ for any sample $i$ if its additional modality were acquired. However, the following upper-bound heuristics are label-agnostic, i.e., the true label ${\textnormal{y}}_{i}$ of a candidate sample is not used when determining its priority for acquisition. Consequently, the selection principle for these strategies must rely on how the known change from an initial score ${\textnormal{s}}_{i}^{\textnormal{avail}}$ to the future score ${\textnormal{s}}_{i}^{\textnormal{full}}$ is expected to influence the global evaluation metric, without direct reference to the sample’s ground-truth label.
@@
The acquisition strategy operates with knowledge of the initial probability $p_{i}^{\textnormal{avail}}=\sigma({\textnormal{s}}_{i}^{\textnormal{avail}})$ derived from the available modalities, and crucially, the true future probability $p_{i}^{\textnormal{full}}=\sigma({\textnormal{s}}_{i}^{\textnormal{full}})$ that would be obtained if the additional modality were acquired (where ${\textnormal{s}}_{i}^{\textnormal{full}}$ is the oracle score). The acquisition score $g_{i}^{\text{UR}}$ for sample $i$ is then the exact reduction in entropy:
$g_{i}^{\text{UR}}=\mathcal{H}(p_{i}^{\textnormal{avail}})-\mathcal{H}(p_{i}^{\textnormal{full}}).$ (14)
@@
$g_{i}^{\text{RC}}=|R(p_{i}^{\textnormal{full}})-R(p_{i}^{\textnormal{avail}})|.$ (15)
@@
The KL-Divergence acquisition strategy aims to identify samples for which acquiring the additional modality would lead to the largest change in the predicted probability distribution. Specifically, it quantifies the divergence from the predicted probability distribution based on the true future score, $P_{i}^{\textnormal{full}}\sim\text{Bernoulli}(p_{i}^{\textnormal{full}})$, back to the initial distribution based on baseline data, $P_{i}^{\textnormal{avail}}\sim\text{Bernoulli}(p_{i}^{\textnormal{avail}})$. This is measured by the KL-Divergence $D_{\mathrm{KL}}(P_{i}^{\textnormal{avail}}\|P_{i}^{\textnormal{full}})$ and can be defined as follows for an acquisition function:
$g_{i}^{\text{KLD}}$ $=D_{\mathrm{KL}}\left(P_{i}^{\textnormal{avail}}\middle\|P_{i}^{\textnormal{full}}\right)$ (16)
$=p_{i}^{\textnormal{avail}}\log_{2}\frac{p_{i}^{\textnormal{avail}}}{p_{i}^{\textnormal{full}}}+(1-p_{i}^{\textnormal{avail}})\log_{2}\frac{1-p_{i}^{\textnormal{avail}}}{1-p_{i}^{\textnormal{full}}}$ (17)
@@ C.4 Baseline Information Strategies
Shifting from approaches that leverage oracle knowledge of future scores (${\textnormal{s}}_{i}^{\textnormal{full}}$), the present section details methods serving as practical, label-agnostic baselines. They make acquisition decisions based exclusively on information derived from the initially available modality (${\textnormal{s}}_{i}^{\textnormal{avail}}$). A random acquisition strategy serves as a fundamental baseline.
@@
The Maximum Baseline Uncertainty strategy is a baseline that prioritizes samples for which the prediction based on the initially available modality is most uncertain. The acquisition score for sample $i$ is directly the binary entropy $\mathcal{H}(p_{i}^{\textnormal{avail}})$, as defined in Equation˜13:
@@ C.5 Imputation-Based Strategies
Having explored strategies that assume perfect knowledge of the true labels ${\textnormal{y}}_{i}$ and/or future scores ${\textnormal{s}}_{i}^{\textnormal{full}}$, and simpler baselines relying only on current information ${\textnormal{s}}_{i}^{\textnormal{avail}}$, we now introduce methods aiming to bridge the gap by offering a practical and label-agnostic pathway to modality acquisition. They operate by utilizing an imputation model, $f_{\text{imp}}$, to generate a set of $K$ plausible future scores, denoted $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$, conditioned on the initially available data ${\textnormal{s}}_{i}^{\textnormal{avail}}$. The core principle of these strategies is to then derive acquisition scores from statistics of this imputed score distribution, with the goal of emulating the decision-making process, but without requiring true future knowledge at test time.
@@
The Maximum Expected Uncertainty Reduction strategy aims to select samples for which the acquisition of the additional modality is anticipated to yield the largest average decrease in predictive uncertainty (Equation˜13). This strategy considers the initial entropy $\mathcal{H}(p_{i}^{\textnormal{avail}})$, and the distribution of entropies $\{\mathcal{H}({\textnormal{p}}_{i,k}^{\textnormal{imp}})\}_{k=1}^{K}$. The acquisition score $g_{i}^{\text{eUR}}$ is the difference between the initial entropy and the mean of the imputed future entropies:
@@ Appendix D Hyperparameters, Model Details and Compute Environment
We employ domain-specific encoders to process the respective modalities: for language inputs, we use a pre-trained BERT model (Devlin et al., 2019), for vision, a Vision Transformer (ViT) (Dosovitskiy et al., 2021). Other data types, e.g., temporal sequences, tabular data, or pre-extracted embeddings, are handled by Transformer encoders (Vaswani et al., 2017). We use well-established hyperparameters from the literature for the modality-specific encoders and only optimize the remaining parameters. Notably, our experiments compared three approaches for normalizing the encoder output: No Normalization, Batch Normalization, and Layer Normalization. We found Layer Normalization to be particularly advantageous, as it both stabilized training convergence and significantly enhanced the performance of the DDPMs. We also evaluated the impact of using only the CLS token representation from the encoder versus leveraging the full output sequence. This comparison revealed no substantial effect on performance, suggesting the sufficiency of the CLS token representation for our task. Layers in the network are initialized using He initialization (He et al., 2015) if they were not pre-initialized by the specific encoder architecture. We find this particularly important for stabilizing the DDPMs during the early epochs of end-to-end model training.
@@
Figure 4: The latent DDPM with its (de)noising functions. Coloring represents less noise in the latent space, starting with pure noise in $X_{i,T}=X_{1,T}$ with $T$ steps and modality ii. The DDPM is conditioned with two non-missing latent spaces, each from one remaining modality respectively.

### v3

@@ C.2 Oracle Acquisition Strategies: Exact Gain Calculation
Oracle acquisition strategies serve as theoretical upper limits for the performance of greedy acquisition approaches. They operate under the ideal assumption that the true labels ${\textnormal{y}}_{i}$ and the outcome scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ are known for all samples $i\in\{1,\dots,N\}$. While not implementable in practice, these oracle strategies provide benchmarks by selecting samples based on their exact marginal contribution to the global evaluation metric. The general principle is to iteratively select $\beta$ samples. At each step, among the samples for which the additional modality has not yet been acquired, the oracle picks the one that provides the largest true immediate gain to the chosen global metric.
@@ C.3 Upper-Bound Heuristic Strategies
The preceding oracle strategies make the assumption of perfect foresight into both the true labels ${\textnormal{y}}_{i}$ and the exact outcome scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$. We now introduce a distinct class of upper-bound heuristic strategies. These strategies still presume access to the true future scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ for any sample $i$ if its additional modality were acquired. However, the following upper-bound heuristics are label-agnostic, i.e., the true label ${\textnormal{y}}_{i}$ of a candidate sample is not used when determining its priority for acquisition. Consequently, the selection principle for these strategies must rely on how the known change from an initial score ${\textnormal{s}}_{i}^{\textnormal{avail}}$ to the future score ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ is expected to influence the global evaluation metric, without direct reference to the sample’s ground-truth label.
@@
The acquisition strategy operates with knowledge of the initial probability $p_{i}^{\textnormal{avail}}=\sigma({\textnormal{s}}_{i}^{\textnormal{avail}})$ derived from the available modalities, and crucially, the true future probability $p_{i}^{\textnormal{acquired}}=\sigma({\textnormal{s}}_{i}^{\textnormal{acquired}})$ that would be obtained if the additional modality were acquired (where ${\textnormal{s}}_{i}^{\textnormal{acquired}}$ is the oracle score). The acquisition score $g_{i}^{\text{UR}}$ for sample $i$ is then the exact reduction in entropy:
$g_{i}^{\text{UR}}=\mathcal{H}(p_{i}^{\textnormal{avail}})-\mathcal{H}(p_{i}^{\textnormal{acquired}}).$ (16)
@@
$g_{i}^{\text{RC}}=|R(p_{i}^{\textnormal{acquired}})-R(p_{i}^{\textnormal{avail}})|.$ (17)
@@
The KL-Divergence acquisition strategy aims to identify samples for which acquiring the additional modality would lead to the largest change in the predicted probability distribution. Specifically, it quantifies the divergence from the predicted probability distribution based on the true future score, $P_{i}^{\textnormal{acquired}}\sim\text{Bernoulli}(p_{i}^{\textnormal{acquired}})$, back to the initial distribution based on baseline data, $P_{i}^{\textnormal{avail}}\sim\text{Bernoulli}(p_{i}^{\textnormal{avail}})$. This is measured by the KL-Divergence $D_{\mathrm{KL}}(P_{i}^{\textnormal{avail}}\|P_{i}^{\textnormal{acquired}})$ and can be defined as follows for an acquisition function:
$g_{i}^{\text{KLD}}$ $=D_{\mathrm{KL}}\left(P_{i}^{\textnormal{avail}}\middle\|P_{i}^{\textnormal{acquired}}\right)$ (18)
$=p_{i}^{\textnormal{avail}}\log_{2}\frac{p_{i}^{\textnormal{avail}}}{p_{i}^{\textnormal{acquired}}}+(1-p_{i}^{\textnormal{avail}})\log_{2}\frac{1-p_{i}^{\textnormal{avail}}}{1-p_{i}^{\textnormal{acquired}}}$ (19)
@@ C.4 Baseline Information Strategies
Shifting from approaches that leverage oracle knowledge of future scores (${\textnormal{s}}_{i}^{\textnormal{acquired}}$), the present section details methods serving as practical, label-agnostic baselines. They make acquisition decisions based exclusively on information derived from the initially available modality (${\textnormal{s}}_{i}^{\textnormal{avail}}$). A random acquisition strategy serves as a fundamental baseline.
@@
The Maximum Baseline Uncertainty strategy is a baseline that prioritizes samples for which the prediction based on the initially available modality is most uncertain. The acquisition score for sample $i$ is directly the binary entropy $\mathcal{H}(p_{i}^{\textnormal{avail}})$, as defined in Equation˜15:
@@ C.5 Imputation-Based Strategies
Having explored strategies that assume perfect knowledge of the true labels ${\textnormal{y}}_{i}$ and/or future scores ${\textnormal{s}}_{i}^{\textnormal{acquired}}$, and simpler baselines relying only on current information ${\textnormal{s}}_{i}^{\textnormal{avail}}$, we now introduce methods aiming to bridge the gap by offering a practical and label-agnostic pathway to modality acquisition. They operate by utilizing an imputation model, $f_{\text{imp}}$, to generate a set of $K$ plausible future scores, denoted $\{{\textnormal{s}}_{i,k}^{\textnormal{imp}}\}_{k=1}^{K}$, conditioned on the initially available data ${\textnormal{s}}_{i}^{\textnormal{avail}}$. The core principle of these strategies is to then derive acquisition scores from statistics of this imputed score distribution, with the goal of emulating the decision-making process, but without requiring true future knowledge at test time.
@@
The Maximum Expected Uncertainty Reduction strategy aims to select samples for which the acquisition of the additional modality is anticipated to yield the largest average decrease in predictive uncertainty (Equation˜15). This strategy considers the initial entropy $\mathcal{H}(p_{i}^{\textnormal{avail}})$, and the distribution of entropies $\{\mathcal{H}({\textnormal{p}}_{i,k}^{\textnormal{imp}})\}_{k=1}^{K}$. The acquisition score $g_{i}^{\text{eUR}}$ is the difference between the initial entropy and the mean of the imputed future entropies:
@@ Appendix D Hyperparameters, Model Details and Compute Environment
We employ domain-specific encoders to process the respective modalities: for language inputs, we use a pre-trained BERT model (devlin_bert_2019), for vision, a Vision Transformer (ViT) (dosovitskiy_image_2021). Other data types, e.g., temporal sequences, tabular data, or pre-extracted embeddings, are handled by Transformer encoders (vaswani_attention_2017). We use well-established hyperparameters from the literature for the modality-specific encoders and only optimize the remaining parameters. Notably, our experiments compared three approaches for normalizing the encoder output: No Normalization, Batch Normalization, and Layer Normalization. We found Layer Normalization to be particularly advantageous, as it both stabilized training convergence and significantly enhanced the performance of the DDPMs. We also evaluated the impact of using only the CLS token representation from the encoder versus leveraging the full output sequence. This comparison revealed no substantial effect on performance, suggesting the sufficiency of the CLS token representation for our task. Layers in the network are initialized using He initialization (he_delving_2015) if they were not pre-initialized by the specific encoder architecture. We find this particularly important for stabilizing the DDPMs during the early epochs of end-to-end model training.
@@
Figure 4: The latent DDPM with its (de)noising functions. Coloring represents less noise in the latent space, starting with pure noise in $X_{i,T}=X_{1,T}$ with $T$ steps. The DDPM is conditioned with two non-missing latent spaces, each from one remaining modality respectively.